# Flavor diagonal nucleon charges using clover fermions on MILC HISQ ensembles

### v2

@@ Abstract
We present lattice results for the flavor diagonal charges of the proton from the analysis of eight ensembles generated using 2+1+1-flavors of highly improved staggered quarks (HISQ) by the MILC collaboration. The calculation includes all the needed connected and disconnected contributions to nucleon three-point function. For extracting matrix elements using fits to the spectral decomposition of these correlation functions, two strategies to remove excited state contributions are employed and compared. To renormalize these charges, the 2+1-flavor mixing matrix is calculated in the RI-sMOM intermediate scheme on the lattice. The final results are presented in the ${\overline{\text{MS}}}$ scheme at scale 2 GeV. The axial charges for the proton are $g_{A}^{u}=0.781(25)$, $g_{A}^{d}=-0.440(39)$, and $g_{A}^{s}=-0.055(9)$; the tensor charges are $g_{T}^{u}=0.782(28)$, $g_{T}^{d}=-0.195(10)$, and $g_{T}^{s}=-0.0016(12)$; and the scalar charges are $g_{S}^{u}=9.39(88)$, $g_{S}^{d}=8.84(93)$, and $g_{S}^{s}=0.37(14)$. Results for the neutron are given by the $u\leftrightarrow d$ interchange. Results for the sigma terms are $\sigma_{\pi N}|_{\rm standard}=42(6)~{}{\rm MeV}$ from a “standard” analysis and $\sigma_{\pi N}|_{N\pi}=61(6)~{}{\rm MeV}$ from a “$N\pi$” analysis that includes the contributions of multihadron $N\pi$ excited states as motivated by chiral perturbation theory. Our preferred value $\sigma_{\pi N}|_{N\pi}$ is consistent with the phenomenological extraction from $\pi-N$ scattering data. The strangeness content of the proton, for which the “standard” analysis is appropriate, is $\sigma_{s}|_{\rm standard}=35(13)~{}{\rm MeV}$.
@@ I Introduction
The motivation for these calculations and much of the methodology used has already been published for $g_{A}^{q}$ in Ref. [2], $g_{T}^{q}$ in [3] and for the pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}\times g_{S}^{u+d}$, in [4]. The charges $g_{A}^{u,d,s}$ give the contributions of the intrinsic spin of the quarks to the nucleon spin; the $g_{T}^{u,d,s}$ give the contribution of the quark electric dipole moment (EDM) operator to the nucleon EDM; and all three, $g_{A,S,T}^{u,d,s}$, give the coupling of dark matter or Higgs-like interactions with nucleons in the respective channels. In addition, with $g_{S}^{u,d,s}$ in hand, we calculate the pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}g_{S}^{u+d}$ and the strangeness content of the nucleon, $\sigma_{s}=m_{s}g_{S}^{s}$. We compare our results with those given in the latest FLAG report 2024 [5], a community review of calculations done until May 2024 by various lattice collaborations, and new results since then in Section VII.
@@
  • Results for $g_{S}^{u,d,s}$ are new. In Ref. [4], the result for only the renormalization group invariant pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}\times g_{S}^{u+d}$, was presented. We use the fact that $m\overline{\psi}\psi$ is the mass term in the action, which implies $Z_{m}\times Z_{S}=1$.
  • Complete results for the renormalization of quark bilinears for the clover-on-HISQ formulation, i.e., the 2+1-flavor mixing matrix, $Z_{A,S,T}$, are obtained nonperturbatively on the lattice using the regularization independent symmetric momentum subtraction (RI-sMOM) scheme [8, 9]. These $Z$s are then matched to the $\overline{MS}$ scheme and run to $2$ GeV using perturbation theory. Results in the basis $(u-d,u+d,s)$, most relevant to this work, are given in Appendix B. Our previous calculations used $Z_{\text{isoscalar}}\approx Z_{\text{isovector}}$, i.e., $Z_{\text{u+d}}\approx Z_{\text{u-d}}$ for the renormalization constants of the axial and tensor operators. Having completed the full calculation, we can now confirm that this is a very good approximation.
@@
    For correlation functions with resolvable signal of ESC, fits to get ground state matrix elements (GSME), we now include two excited states in the spectral decomposition, and in each case compare fits with and without including the $N\pi$ excited state. These fits have been made using the full covariance matrix. To further control ESC, we vary $t_{\rm skip}$, the number of points skipped adjacent to the source and the sink, and the range of $\tau$ values included in the fits. For final results we take the average of these various fits, weighted by the Akaika Information Criteria (AIC) score [13].
@@
Table 1: The lattice spacing $a$, the valence pion mass $M_{\pi}$, gauge coupling $\beta$, the Sheikholeslami-Wohlert coefficient $C_{SW}$ defining the clover term in the action, the light quark mass $am_{ud}=1/2\kappa_{ud}-4$, the strange quark mass $am_{s}=1/2\kappa_{s}-4$, the lattice volume $L^{3}\times T$, and the lattice size in units of $M_{\pi}$ for the eight ensembles analyzed for flavor diagonal charges $g_{\Gamma}^{q}$.
@@
$a09m130$ 7.0 {10,12,14,16} 1274 1290 128 4 1270 10000 50 994 $(^{\dagger})$ 50
$a06m310$ 12.0 {18,20,22,24} 977 500 128 4 808 12000 50 976 $(^{\dagger})$ 50
$a06m220$ 11.0/9.0 (*) {18,20,22,24} 1010 649 64 4 1001 10000 50 1002 10000 50
Table 2: Parameters used in the calculation of the charges $g_{\Gamma}^{q}$ on the eight ensembles. The second column gives the smearing parameter $\sigma$ used in the creation of the covariant Gaussian smeared source for quark propagators. For ensemble $a06m220$, marked with (*), the connected contribution was done with $\sigma=11$ while the disconnected was done with $\sigma=9$. The list of source-sink time separations $\tau$ analyzed for the connected 3-point functions is given in column three. The disconnected contributions were calculated for all $\tau$ except on $a12m220$ marked with (**). $N_{\text{conf}}^{\rm 2pt}$ and $N_{\text{conf}}^{\rm conn}$ give the number of configurations analyzed for 2-point and connected 3-point correlation functions, respectively. In the truncated solver with bias correction method, we used $N_{\text{LP}}$ low and $N_{\text{HP}}$ high precision measurements per configuration for the connected contributions. In columns 8–10 (11–13), $N_{\text{conf}}^{l}$ ($N_{\text{conf}}^{s}$) configurations were analyzed for the light (strange) quark disconnected contributions with $N_{\rm src}^{l}$ ($N_{\rm src}^{s}$) $Z_{4}$ random noise sources and $N_{\rm LP}/N_{\rm HP}$ low to high precision measurements. In cases marked with $(^{\dagger})$, the $N^{s}_{\text{conf}}$ configurations were split into two sets and analyzed with different number of sources $N_{\rm src}^{s}$, and the results averaged. The analysis of the isovector charges on the full set of thirteen configurations has been published in Ref. [14].
@@ II Lattice Methodology
where $\mathcal{P}$ is the spin projection defined in Ref. [15]. The interpolating operator ${\mathcal{N}}$ used to create/annihilate the nucleon at the source/sink time slices 00 and $\tau$ is
@@
To reduce the computational cost, the connected diagrams are constructed using a coherent sequential source [20, 18] and the truncated solver with bias correction [21, 22, 18] methods. The quark loop with zero-momentum operator insertion, i.e., $\mathbf{q=0}$, for the disconnected contribution, $C_{\Gamma}^{\text{disc}}$, is estimated stochastically using $Z_{4}$ random noise sources. This calculation is accelerated with a combination of the truncated solver with bias correction method and the hopping parameter expansion [23, 24]. For the scalar case, the disconnected contribution is calculated using the vacuum subtracted operator $\mathcal{O}_{S}^{q}-\langle\mathcal{O}_{S}^{q}\rangle$ [25]. Additional details are given in Ref. [15]. The statistics for the connected and disconnected calculations are given in Table 2.
@@
In our previous works [2, 3], the removal of excited-state contamination (ESC) was carried out by making separate fits to $C^{\text{conn}}_{\Gamma}(t;\tau)$ and $C^{\text{disc}}_{\Gamma}(t;\tau)$. Similarly, the chiral-continuum (CC) extrapolations of $g_{\Gamma}^{q,\text{disc}}$ and $g_{\Gamma}^{q,\text{conn}}$ to $a\to 0$ and $M_{\pi}\to 135$ MeV were performed independently. This approach may have introduced an unquantified systematic bias as discussed in Ref. [2]. This systematic has been removed in this work by making ESC fits to the sum and the subsequent CCFV extrapolation of the renormalized charges $g_{\Gamma}^{q}$. To understand the size of possible systematics in our previous work, we obtained results (ESC fits and CCFV extrapolation) both ways, and find the differences to be smaller than the statistical errors.
@@ III Strategies to Remove Excited-State Contributions
The challenge lies in removing all ESC since, even in our highest statistics ensembles, the right hand side of Eq. (5) is truncated at 3-states. While the ESC decrease exponentially with source-sink separation $\tau$, the signal-to-noise ratio in the $C^{\textrm{3pt}}_{\Gamma}$ also degrades exponentially as $e^{-(M_{N}-3/2M_{\pi})\tau}$ [34, 35]. The $C^{\textrm{3pt}}_{\Gamma}$ are well-measured only up to source-sink separation $\tau\approx 1.4\mathop{\rm fm}\nolimits$ and for $\tau\lesssim 1.4\mathop{\rm fm}\nolimits$, ESC are found to be significant. Three- state fits to remove these using Eq. (1) with $\mathcal{A}_{0}$, $M_{i}$ and the $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ left as free parameters are not stable with current statistics. Additional information on these parameters is needed.
To make robust fits to Eq. (5), the key parameters are the ground state amplitude $\mathcal{A}_{0}$ and the finite volume spectrum $M_{i}$. If these are known a priori, then only the combinations $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ involving the excited states remain as free parameters to extract $\langle 0|O_{\Gamma}^{q}|0\rangle$. This then becomes a tractable problem. Note, these $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ are not used in any subsequent analysis.
@@
To make n-state fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ with input from $C^{\textrm{2pt}}(\tau)$, we need $C^{\textrm{2pt}}(\tau)$ truncated to $n+1$ (or higher) number of states presuming that the effects of all the neglected states are subsumed in the parameters of the $n+1$ (and higher) state. Of these, $M_{1}$ is the most important for controlling ESC in fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ as the contamination falls exponentially with the mass gap. Unfortunately, with current statistics, 4-state fits to Eq. (6) give large regions of parameter space with roughly the same $\chi^{2}/dof$. We, therefore, resort to using physics motivated priors.
@@
The second challenge is that fits to $C^{\textrm{2pt}}(\tau)$ do not expose multihadron (such as $N\pi$, $N\pi\pi,\ldots$) excited states. The reason is that the coupling of multihadron states with nucleon quantum numbers to $\cal{N}$ is suppressed by factors of $1/L^{3}$ for each extra state [36], however, as discussed later, their contribution to the matrix elements of certain operators is enhanced and compensates for the volume suppression. Therefore, in cases of enhanced contributions, these states need to be included in fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ because the mass of $N(1)\pi(-1)$ (or $N(0)\pi(0)\pi(0)$) state is about 1230 MeV, which is significantly smaller than the first radial excitation $N(1440)$. In short, our first goal is to determine all the excited states, i.e., their $M_{i}$, that make significant contributions.
@@
Our overall approach consists of making simultaneous fits to $C^{\textrm{2pt}}(\tau)$ (truncated at four states) and $C^{\textrm{3pt}}_{\Gamma}(t;\tau)$ (truncated at 2 or 3 states) within a single elimination jackknife process and using the full covariance matrix. In fits without priors, or wide priors, the final $A_{i}$ and $M_{i}$ are predominantly fixed by $C^{\textrm{2pt}}(\tau)$ and the $C^{\textrm{3pt}}_{\Gamma}(t;\tau)$ give $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ including the GSME $\langle 0|O_{\Gamma}^{q}|0\rangle$. In this case the priors are used mainly to stabilize the fits. The drawback, with current statistics, is that there is a whole region in the $\{A_{i},M_{i}\}$ space that gives fits to $C^{\textrm{2pt}}(\tau)$ with similar $\chi^{2}/dof$. Thus, there is considerable range of plausible values for $M_{1}$ with the lowest, theoretically, being the $N(1)\pi(-1)$ or $N(0)\pi(0)\pi(0)$. To reduce this uncertainty one can input physically motivated values of the $M_{i}$ through narrow priors.
@@
Ideally, one wants a data driven analysis, i.e., the data and the fit should fix the $M_{i}$. With the current statistics, we are not able to achieve this desired situation. We, therefore, carry out the full analysis using two physically motivated values for $M_{1}$:
@@
    The “standard” fit to $C^{\textrm{2pt}}(\tau)$ and $C^{\textrm{3pt}}(t;\tau)$ employs wide priors for all excited-state amplitudes, $\mathcal{A}_{i}$, and masses, $M_{i}$. In this case the priors are used only to stabilize the fits. These $A_{0}$ and $M_{i}$, are mainly constrained by $C^{\textrm{2pt}}(\tau)$. The value of $M_{1}$ found in this case is $\gtrsim 1.5$ GeV. This simultaneous fit is labeled “$\{4\}$” and $M_{1}$ is loosely identified with the radial excitation $N(1440)$.
@@
    The “$N\pi$” fit uses a narrow prior for $M_{1}$, centered on the non-interacting energy of the lowest allowed $N\pi$ (or $N\pi\pi$) lattice state on that ensemble. In practice, the energy of these two states is very similar, so they can be considered as one in Eq. (6) with the joint amplitude given by the sum. Throughout this work, for brevity we use the label $N\pi$ for their combined contribution. The priors for the third and fourth state are the same as in the “standard” fit. This fit is labeled “$\{4^{N\pi}\}$”.
Further details of our implementation of these two fits strategies are given in Refs. [11, 4]. We carry out a full analysis of the nucleon charges with these two choices of $M_{1}$. Since the two fit strategies are not distinguished by the $\chi^{2}$, we explain, when quoting the final results, the physics motivations for the choices made. Also, where appropriate, we quote the difference as an estimate of an associated systematic uncertainty.
It is important to note that the mass of the $N\pi$ or the $N\pi\pi$ state decreases below that for the first radial excitation $N(1440)$ only for ensembles with $M_{\pi}<200$ MeV, reaching $\sim 1230$ MeV for physical pion mass ensembles. Thus, in cases with enhanced matrix elements, one expects to observe a strong dependence on $M_{\pi}$ only for $M_{\pi}<200$ MeV, which in our data leaves only the $a09m130$ ensemble. Fits to these data are shown in Figs. 2 and 3.

### v3

@@ Abstract
We present lattice results for the flavor diagonal charges of the proton from the analysis of eight ensembles generated using 2+1+1-flavors of highly improved staggered quarks (HISQ) by the MILC collaboration. The calculation includes all the needed connected and disconnected contributions to nucleon three-point function. For extracting matrix elements using fits to the spectral decomposition of these correlation functions, two strategies to remove excited state contributions are employed and compared. To renormalize these charges, the 2+1-flavor mixing matrix is calculated in the RI-sMOM intermediate scheme on the lattice. The final results are presented in the $\overline{\text{MS}}$} scheme at scale 2 GeV. The axial charges for the proton are $g_{A}^{u}=0.781(25)$, $g_{A}^{d}=-0.440(39)$, and $g_{A}^{s}=-0.055(9)$; the tensor charges are $g_{T}^{u}=0.782(28)$, $g_{T}^{d}=-0.195(16)$, and $g_{T}^{s}=-0.0016(12)$; and the scalar charges are $g_{S}^{u}=9.39(88)$, $g_{S}^{d}=8.84(93)$, and $g_{S}^{s}=0.37(14)$. Results for the neutron are given by the $u\leftrightarrow d$ interchange. Results for the sigma terms are $\sigma_{\pi N}|_{\rm standard}=42(6)~{\rm MeV}$ from a “standard” analysis and $\sigma_{\pi N}|_{N\pi}=61(6)~{\rm MeV}$ from a $N\pi$” analysis that includes the contributions of multihadron $N\pi$ excited states as motivated by chiral perturbation theory. Our preferred value $\sigma_{\pi N}|_{N\pi}$ is consistent with the phenomenological extraction from $\pi-N$ scattering data. The strangeness content of the proton, for which the “standard” analysis is appropriate, is $\sigma_{s}|_{\rm standard}=35(13)~{\rm MeV}$.
@@ I Introduction
The motivation for these calculations and much of the methodology used has already been published for $g_{A}^{q}$ in Ref. [2], $g_{T}^{q}$ in [3] and for the pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}\times g_{S}^{u+d}$, in [4]. The charges $g_{A}^{u,d,s}$ give the contributions of the intrinsic spin of the quarks to the nucleon spin; the $g_{T}^{u,d,s}$ give the contribution of the quark electric dipole moment (EDM) operator to the nucleon EDM; and all three, $g_{A,S,T}^{u,d,s}$, give the coupling of dark matter or Higgs-like interactions with nucleons in the respective Lorentz channels. In addition, with $g_{S}^{u,d,s}$ in hand, we calculate the pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}g_{S}^{u+d}$ and the strangeness content of the nucleon, $\sigma_{s}=m_{s}g_{S}^{s}$. We compare our results with those given in the latest FLAG report 2024 [5], a community review of calculations done until May 2024 by various lattice collaborations, and new results since then in Section VII.
@@
  • Results for $g_{S}^{u,d,s}$ are new. In Ref. [4], the result for only the renormalization group invariant pion-nucleon sigma term, $\sigma_{\pi N}=m_{u,d}\times g_{S}^{u+d}$, was presented. Using the fact that $m\overline{\psi}\psi$ is the mass term in the action, implies $Z_{m}\times Z_{S}=1$.
  • Complete results for the renormalization of quark bilinears for the clover-on-HISQ formulation, i.e., the 2+1-flavor mixing matrix, $Z_{A,S,T}$, are obtained nonperturbatively on the lattice using the regularization independent symmetric momentum subtraction (RI-sMOM) scheme [8, 9]. These $Z$s are then matched to the $\overline{MS}$ scheme and run to $2$ GeV using results from perturbation theory. Results in the basis (ud,u+d,s)(u-d,u+d,s), most relevant to this work, are given in Appendix B. Our previous calculations used $Z_{\text{isoscalar}}\approx Z_{\text{isovector}}$, i.e., $Z_{\text{u+d}}\approx Z_{\text{u-d}}$ for the renormalization constants of the axial and tensor operators. Having completed the full calculation, we find removing this approximation has not changed the results significantly because the corrections are small as can be inferred from the numbers in Appendix B.
@@
    For correlation functions with resolvable signal of ESC, fits to get ground state matrix elements (GSME) now include two excited states in the spectral decomposition, and in each case compare fits with and without including the $N\pi$ excited state. These fits have been made using the full covariance matrix. To further control ESC, we vary $t_{\rm skip}$, the number of points skipped adjacent to the source and the sink, and the range of $\tau$ values included in the fits. For final results we take the average of these various fits, weighted by the Akaika Information Criteria (AIC) score [13].
@@
Table 1: The lattice spacing $a$, the valence pion mass $M_{\pi}$, gauge coupling $\beta$, the Sheikholeslami-Wohlert coefficient $C_{SW}$ defining the clover term in the Wilson action, the light quark mass $am_{ud}=1/2\kappa_{ud}-4$, the strange quark mass $am_{s}=1/2\kappa_{s}-4$, the lattice volume $L^{3}\times T$, and the lattice size in units of $M_{\pi}$ for the eight ensembles analyzed for flavor diagonal charges $g_{\Gamma}^{q}$.
@@
$a09m130$ 7.0 {10,12,14,16} 1274 1290 128 4 1270 10000 50 994 10000/4000()($^{\dagger}$) 50
$a06m310$ 12.0 {18,20,22,24} 977 500 128 4 808 12000 50 976 10000/4000()($^{\dagger}$) 50
$a06m220$ 11.0/9.0 (*) {18,20,22,24} $10$ 649 64 4 1001 10000 50 1002 10000 50
Table 2: Parameters used in the calculation of the charges $g_{\Gamma}^{q}$ on the eight ensembles. The second column gives the smearing parameter $\sigma$ used in the creation of the covariant Gaussian smeared source for quark propagators. For ensemble $a06m220$, marked with (*), the connected contribution was done with $\sigma=11$ while the disconnected was done with $\sigma=9$. The list of source-sink time separations $\tau$ analyzed for the connected 3-point functions is given in column three. The disconnected contributions were calculated for all $\tau$ and all intermediate $t$ except on $a12m220$ for light quarks ($N^{l}_{\rm conf}$ marked with (#)). On this ensemble, data were collected with fixed $t_{\rm skip}=3$ and the four source time slices were fixed at $t_{src}=2,18,34,50$ rather than being chosen randomly. $N_{\text{conf}}^{\rm 2pt}$ and $N_{\text{conf}}^{\rm conn}$ give the number of configurations analyzed for 2-point and connected 3-point correlation functions, respectively. In the truncated solver with bias correction method, we used $N_{\text{LP}}$ low and $N_{\text{HP}}$ high precision measurements per configuration for the connected contributions. Columns 8–10 (11–13), specify the number of configurations, $N_{\text{conf}}^{l}$ ($N_{\text{conf}}^{s}$), analyzed for the light (strange) quark disconnected contributions with $N_{\rm src}^{l}$ ($N_{\rm src}^{s}$) $Z_{4}$ random noise sources and $N_{\rm LP}/N_{\rm HP}$ low to high precision measurements. In cases marked with $(^{\dagger})$, the $N^{s}_{\text{conf}}$ configurations were split into two sets and analyzed with different number of sources $N_{\rm src}^{s}$, and the results averaged. Results for isovector charges on full set of thirteen configurations including these has been published in Ref. [14].
@@ II Lattice Methodology
where $\mathcal{P}$ is the spin projection defined in Ref. [15]. The interpolating operator $\mathcal{N}$} used to create/annihilate the nucleon at the source/sink time slices 0 and $\tau$ is
@@
To reduce the computational cost, the connected diagrams are constructed using a coherent sequential source [20, 18] and the truncated solver with bias correction [21, 22, 18] methods. The quark loop with zero-momentum operator insertion, i.e., $\mathbf{q=0}$, for the disconnected contribution, $C_{\Gamma}^{\text{disc}}$, is estimated stochastically using $Z_{4}$ random noise sources. This calculation was accelerated with a combination of the truncated solver with bias correction method [15], and the hopping parameter expansion used as a pre-conditioner [23, 24, 15]. For the scalar case, the disconnected contribution is calculated using the vacuum subtracted operator $\mathcal{O}_{S}^{q}-\langle\mathcal{O}_{S}^{q}\rangle$ [25]. Additional details are given in Ref. [15]. The statistics for the connected and disconnected calculations are given in Table 2.
@@
In our previous works [2, 3], the removal of excited-state contamination (ESC) was carried out by making separate fits to $C^{\text{conn}}_{\Gamma}(t;\tau)$ and $C^{\text{disc}}_{\Gamma}(t;\tau)$. Similarly, the chiral-continuum (CC) extrapolations of $g_{\Gamma}^{q,\text{disc}}$ and $g_{\Gamma}^{q,\text{conn}}$ to $a\to 0$ and $M_{\pi}\to 135$ MeV were performed independently. This procedure may have introduced an unquantified systematic bias as discussed in Ref. [2]. This systematic has been removed in this work by making ESC fits to the sum and the subsequent CCFV extrapolation of the renormalized charges $g_{\Gamma}^{q}$. To understand the size of possible systematics in our previous work, we obtained results (ESC fits and CCFV extrapolation) both ways, and find the differences to be smaller than the statistical errors.
@@ III Strategies to Remove Excited-State Contributions
The challenge lies in removing all ESC since, even in our highest statistics ensembles, the right hand side of Eq. (5) is truncated at 3-states. While the ESC decrease exponentially with the mass gap $\Delta M$ and source-sink separation $\tau$ as $e^{-\Delta M\tau}$, the signal-to-noise ratio for $C^{\textrm{3pt}}_{\Gamma}$ degrades even faster exponentially as $e^{-(M_{N}-3/2M_{\pi})\tau}$ [34, 35]. The $C^{\textrm{3pt}}_{\Gamma}$ are well-measured only up to source-sink separation $\tau\approx 1.4\mathop{\rm fm}\nolimits$ and for $\tau\lesssim 1.4\mathop{\rm fm}\nolimits$, ESC are found to be significant. Three- state fits to remove these using Eq. (5) with $\mathcal{A}_{0}$, $M_{i}$ and the $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ left as free parameters are not stable with current statistics. Additional information on these parameters is needed.
To make robust fits to Eq. (5), the key parameters are the ground state amplitude $\mathcal{A}_{0}$ and the finite volume spectrum $M_{i}$. If these are known a priori, then to extract $\langle 0|O_{\Gamma}^{q}|0\rangle$ only the combinations $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ involving the excited states remain as free parameters. This then becomes a tractable problem. Note, these $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ are not used in any subsequent analysis.
@@
To make n-state fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ with input from $C^{\textrm{2pt}}(\tau)$, we need $C^{\textrm{2pt}}(\tau)$ truncated to $n+1$ (or higher) number of states presuming that the effects of all the neglected states are subsumed in the parameters of the $n+1$ (and higher) state. Of these, $M_{1}$ is the most important for controlling ESC in fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ with the contamination falling exponentially with the mass gap. Unfortunately, with current statistics, 4-state fits to Eq. (6) give large regions of parameter space with roughly the same $\chi^{2}/dof$. We, therefore, resort to using physics motivated priors to determine the $M_{i}$.
@@
The second challenge is that fits to $C^{\textrm{2pt}}(\tau)$ do not expose multihadron (such as $N\pi$, $N\pi\pi,\ldots$) excited states. The reason is that the coupling of multihadron states with nucleon quantum numbers to $\cal{N}$ is suppressed by factors of $1/L^{3}$ for each extra state [36]. On the other hand, as discussed later, their contribution to the matrix elements of certain operators is enhanced and compensates for the volume suppression. Therefore, in cases with enhanced contributions, these states need to be included in fits to $C^{\text{3pt}}_{\Gamma}(t;\tau)$ because the mass of $N(1)\pi(-1)$ (or $N(0)\pi(0)\pi(0)$) state is about 1230 MeV, which is significantly smaller than the first radial excitation $N(1440)$. In short, our first goal is to determine all the excited states, i.e., their $M_{i}$, that make significant contributions.
Our overall approach consists of making simultaneous fits to $C^{\textrm{2pt}}(\tau)$ (truncated at four states) and $C^{\textrm{3pt}}_{\Gamma}(t;\tau)$ (truncated at 2 or 3 states) within a single elimination jackknife process using binned data. [^1] The binning of the data is done in two steps. First, the $O(100)$ measurements made on each configuration are averaged. Second, these configuration averages are further binned using bin sizes ranging between 2–9 resulting in 250–325 data points for the various ensembles. The auto-correlation function calculated using these binned data, which enter in the single elimination jackknife process, show no measurable correlations. Consequently, no augmentation of the single elimination jackknife errors is deemed necessary. All fits are made using the full covariance matrix with respect to the time slices included in the fits. It is, however, restricted to being block diagonal for simultaneous fits to the two-and three-point correlation functions as the off diagonal elements between them are not well-determined with the current statistics.
We also carried out separate fits to the two- and three-point functions and the two sets of results overlap. No anomalous case showing a bias in the simultaneous fits was found. Since the simultaneous fits give more conservative errors and are better motivated as a method, the results are presented using them.
@@
In fits without priors, or wide priors, the $A_{i}$ and $M_{i}$ are predominantly fixed by $C^{\textrm{2pt}}(\tau)$, while the $C^{\textrm{3pt}}_{\Gamma}(t;\tau)$ give $A_{i}^{\ast}A_{j}\langle i|O_{\Gamma}^{q}|j\rangle$ including the GSME $\langle 0|O_{\Gamma}^{q}|0\rangle$. In this case the priors are used mainly to stabilize the fits. The drawback, with current statistics, is that there is a whole region in the $\{A_{i},M_{i}\}$ space that gives fits to $C^{\textrm{2pt}}(\tau)$ with similar $\chi^{2}/dof$. Thus, there is considerable range of plausible values for $M_{1}$ with the lowest, theoretically, being the $N(1)\pi(-1)$ or $N(0)\pi(0)\pi(0)$. One way to reduce this uncertainty is to input physically motivated values of the $M_{i}$ through narrow priors.
@@
Ideally, one wants a data driven analysis procedure, i.e., the data and the fit should fix the $M_{i}$. With the current statistics, we are not able to achieve this desired situation. We, therefore, carry out the full analysis using two physically motivated but different values for $M_{1}$:
@@
    The “standard” fit to $C^{\textrm{2pt}}(\tau)$ and $C^{\textrm{3pt}}(t;\tau)$ employs wide priors for all excited-state amplitudes, $\mathcal{A}_{i}$, and masses, $M_{i}$. In these simultaneous fits, the priors are used only to stabilize the fits. The $A_{0}$ and $M_{i}$, are mainly constrained by $C^{\textrm{2pt}}(\tau)$. The value of $M_{1}$ found in this case is $\gtrsim 1.5$ GeV. This simultaneous fit is labeled $\{4\}$” and $M_{1}$ is loosely identified with the radial excitation $N(1440)$.
@@
    The $N\pi$” fit uses a narrow prior for $M_{1}$, centered on the non-interacting energy of the lowest allowed, $N(1)\pi(-1)$ or $N(0)\pi(0)\pi(0)$, lattice state on that ensemble. In practice, the energies of these two states are very similar, so they can be considered as one in Eq. (6) with a combined amplitude given by the sum. Throughout this work, for brevity we use the label $N\pi$ for their combined contribution. The priors for the other states are the same as in the “standard” fit. This fit is labeled $\{4^{N\pi}\}$”.
Since the two fit strategies are not distinguished by the $\chi^{2}$, we explain, when quoting the final results, the physics motivations for the choices made. Also, where appropriate, we quote the difference as an estimate of an associated systematic uncertainty. Further details of our implementation of these two fits strategies are given in Refs. [11, 4].
It is important to note that the mass of the $N\pi$ or the $N\pi\pi$ state decreases below that for the first radial excitation $N(1440)$ only for ensembles with $M_{\pi}\lesssim 200$ MeV, reaching $\sim 1230$ MeV for physical pion mass ensembles. Thus, in cases with enhanced matrix elements, one expects to observe a strong dependence on $M_{\pi}$ only for $M_{\pi}<200$ MeV, which in our calculation leaves only the $a09m130$ ensemble. Fits to these data are shown in Figs. 2 and 3.